Grade 12Organic Chemistry

What isto be in conjugation with the benzene ring and hence the shared pair of electrons are not available for sharing

In amines i came across this and can u explain me what it is to be in conjugation with a benzene ring ???

Pls explain

Thank a lot !!!

Geoffrey Richards
15 Years agoGrade 12

1 Answer

vikas askiitian expert
ApprovedApproved Tutor Answer15 Years ago

eg=>CH3-CH2-CH=CH-NH2

from ryt side ,  C1-C2 are in conjugation , C1-NH2 are in conjugaion ...

any two atoms are in conjugation if they are having (lone pairs , charge , single electron , double or triple bond) ...

eg=>CH2=CH-Cl

here C-Cl are in conjugation coz Cl has lone paires ..

 

group like Cl , NH2 when attached to benzene ring then these are said to be in conjugation with benzene 

coz benzene contains double bonds & these groups are in conjugation with nearby doubly bonded carbon of ring...